A mixture of CS2(g) and excess O2(g) is placed in a 10 L reaction vessel at 100.0 ∘C and a pressure of 3.10 atm . A spark causes the CS2 to ignite, burning it completely, according to the equation:
CS2(g)+3O2(g)→CO2(g)+2SO2(g)
After reaction, the temperature returns to 100.0 ∘C, and the mixture of product gases (CO2, SO2, and unreacted O2) is found to have a pressure of 2.45 atm .<P 9 Problem 9.109: Chapter Problem Peri < 14 of 20 > Review Constants Periodic Table A mixture of CS2 (9) and excess 02 (9)
Answer #1

V= 10L T = 100.0 C = 373.15 K P1 = 3.10 atm P2 = 2.45 atm m=moles Before Rxn: (n1 + n2) = PV/RT (n1 + n2) = (3.10 atm (10 L)/